# Euclid Quick Data Release (Q1)

### v1

@@ Appendix B Recovery of Einstein mass
We conjecture that the correct recovery could be due to the fact that the network is learning known or unknown strong correlations between some lensing features and the Einstein mass values, bypassing the need for direct inference of redshift values. Some of these features could be lensing distortion patterns (larger Einstein masses produce more pronounced arcs), size and spread of the lensed images (larger Einstein masses produce wider separations of the lensed images), and brightness of the arcs (larger Einstein masses focus light more intensely, which affects the intensity profiles of the arcs in the image). The network could also be learning implicit correlations between features related to the lens redshift, such as the relative brightness of source and lens, and the redshift of the lens itself. The network could then use these correlations to create a Bayesian mapping that reproduces the relation (12), by learning a distribution of possible mappings between Einstein radius and Einstein mass. Thus, when making predictions, the network sees a particular arc size that often correlates with a higher Einstein mass, and gives a value that peaks around the higher value, with smaller or larger uncertainty associated to how often the value of $R_{\textrm{Ein}}$ is associated to a certain value of $M_{\textrm{Ein}}$ in the training phase.
@@
Figure 16 shows the recovery of the Einstein mass for the SLACS Euclidised lenses sample, which is the only one that has values of Einstein mass reported (Auger et al. 2009, table 4). The recovery of the parameter shows a systematic overestimation, with a mean bias of $\mu=0.23$. This is similar, but slightly larger, to the bias for the recovery of the Einstein radius for the Euclidised lenses.
@@
To further investigate the origin of this systematic effect, in Fig. 17 we show the correlation between the Einstein mass and the Einstein radius. For a fixed redshift, we expect ideally a single power-law curve, corresponding to Eq. (12) with fixed values of $D_{\textrm{ds}}$, $D_{\textrm{s}}$ and $D_{\textrm{d}}$. For different redshifts, in log space we expect a family of curves with fixed slope and varying $y$-intercept, which is what we see in the figure with the training set region. The figure shows that the region of the training set corresponding to the minimum and maximum redshifts of the SLACS sample aligns with the lower boundary of the training set area, where most of the SLACS sample is concentrated. The ERO lenses from Euclid, instead, lie fully within the training-set parameter space, which could explain the better recovery of the Einstein radius for these lenses.
Figure 15: Same as Fig. 4, but for the Einstein mass parameter.
Figure 16: Recovery of the Einstein mass parameter for the Euclidised SLACS lenses, showing how well the predicted parameter from LEMON ($y$ coordinate) reproduces the corresponding true value ($x$ coordinate), obtained from Auger et al. (2009).
Figure 17: Einstein mass as a function of Einstein radius. Grey points are the values for the training set, with black points corresponding to those lenses of the training sample having values of lens redshift within the redshift range of the SLACS sample. Orange squares are the literature parameter values for SLACS lenses, while green triangles are the SLACS predicted values from LEMON. Blue points correspond to the predicted values from LEMON for Euclid ERO lenses considered in Fig. 12.

### v2

@@ Appendix B Recovery of Einstein mass
Figure 15: Same as Fig. 4, but for the Einstein mass parameter.
Figure 16: Comparison between predictions of Einstein mass obtained from LEMON and from classical modelling (Auger et al. 2009) for the Euclidised SLACS lenses.
Figure 17: Einstein mass as a function of Einstein radius. Grey points are the values for the training set, with black points corresponding to those lenses of the training sample having values of lens redshift within the redshift range of the SLACS sample. Orange squares are the literature parameter values for SLACS lenses, while green triangles are the SLACS predicted values from LEMON. Red points correspond to the predicted values from LEMON for Euclid ERO lenses considered in Fig. 11.
@@
We conjecture that the correct recovery could be due to the fact that the network is learning known or unknown strong correlations between some lensing features and the Einstein mass values, bypassing the need for direct inference of redshift values. Some of these features could be lensing distortion patterns (larger Einstein masses produce more pronounced arcs), the size and spread of the lensed images (larger Einstein masses produce wider separations of the lensed images), and the brightness of the arcs (larger Einstein masses focus light more intensely, which affects the intensity profiles of the arcs in the image). The network could also be learning implicit correlations between features related to the lens redshift, such as the relative brightness of source and lens, and the redshift of the lens itself. The network could then use these correlations to create a Bayesian mapping that reproduces the relation (12), by learning a distribution of possible mappings between Einstein radius and Einstein mass. Thus, when making predictions, the network sees a particular arc size that often correlates with a higher Einstein mass, and gives a value that peaks around the higher value, with smaller or larger uncertainty associated with how often the value of $R_{\textrm{Ein}}$ is associated with a certain value of $M_{\textrm{Ein}}$ in the training phase.
@@
Figure 16 shows the recovery of the Einstein mass for the SLACS Euclidised lenses sample, which is the only one that has corresponding values from classical modelling of the Einstein mass reported (Auger et al. 2009, table 4). We preventively removed all lenses with effective radii higher than 2, as performed in Sect. 6.1. The comparison between methods shows a systematic overestimation with respect to the classical modelling, with a mean bias of $\mu=10.9$. The negative coefficient of determination also shows a complete mismatch with the results from Auger et al. (2009).
@@
To further investigate this mismatch, in Fig. 17 we show the correlation between the Einstein mass and the Einstein radius. For a fixed redshift, we expect ideally a single power-law curve, corresponding to Eq. (12) with fixed values of $D_{\textrm{ds}}$, $D_{\textrm{s}}$, and $D_{\textrm{d}}$. For different redshifts, in log space we expect a family of curves with fixed slope and varying $y$ intercept, which is what we see in the figure with the training set region. The figure shows that the region of the training set corresponding to the minimum and maximum redshifts of the SLACS sample aligns with the lower boundary of the training set area, where most of the SLACS sample is concentrated. The ERO lenses from Euclid, instead, lie fully within the training-set parameter space, which could explain the better recovery of the Einstein radius for these lenses. We thus conjecture that the mismatch with literature results for this parameter could be due to the fact that the sample lies on the edge of the training set, or that the parameter is intrinsically more difficult to recover with machine learning for real lens images, which are naturally more complex. Future estimates of Einstein mass for Euclid lenses obtained from classical modelling could provide more substantial information on the recovery of this parameter.
Appendix C Robustness of Platt-scaling
Figure 18: Empirical reliability plot generated from the Euclidised lenses’ predictions of Sect. 6.1.
Figure 19: Empirical reliability plot generated from the Q1 lenses’ predictions of Sect. 6.3. Continuous lines are associated with the default calibration obtained from the simulations, while dashed lines are the results after the manual correction with the values from Table 4.
Table 4: Scaling factors needed to produce Gaussian uncertainties for Q1 lenses’ predictions.
Parameter Scaling factor
Einstein radius $R_{\textrm{Ein}}/\textrm{arcsec}$ $1.2$
Mass $x$-ellipticity $\epsilon_{x,\mathrm{m}}$ $1.5$
Mass $y$-ellipticity $\epsilon_{y,\mathrm{m}}$ $1.5$
Light $x$-ellipticity $\epsilon_{x,\mathrm{l}}$ $1.2$
Light $y$-ellipticity $\epsilon_{y,\mathrm{l}}$ $1.2$
Lens effective radius $R_{\textrm{e, lens}}/\textrm{arcsec}$ $2.0$
Lens Srsic index $n_{\textrm{lens}}$ $2.0$
Lens magnitude $m_{\textrm{lens}}$ $1.5$
In this section we discuss the robustness of the Platt-scaling correction for the total uncertainty of our BNN. Due to distributional shift and other systematic effects that produce a difference between simulated and real images, it is possible that the total calibrated uncertainties predicted in output by LEMON, which have been made Gaussian for predictions on the mock Euclid lenses with the Platt-scaling calibration procedure, may become unreliable on real lenses. To check this, we evaluated the empirical reliability plots for both the Euclidised and the Q1 sample of lenses, by using the results (predicted best values and calibrated uncertainties) reported in Sects. 6.1 and 6.3. We show the one for the Euclidised lenses in Fig. 18, while the one for the Q1 lenses in Fig. 19 (continuous lines).
For the Euclidised lenses, the Einstein radius and ellipticity uncertainties are still consistent with the assumption of Gaussianity, as can be seen by the fact that the reliability plot trends for these parameters roughly follow the identity line. The effective radius predictions, instead, show signs of overconfidence, with uncertainties lower than what they should be if they were calibrated correctly. This means that the output uncertainties of our network for this parameter, using the scaling factors that we assumed for the simulated lenses, underestimate the expected uncertainty under the assumption of Gaussianity. This systematic underestimation of the uncertainty appears also for all the predictions of the Q1 lenses.
It should be noted, however, that for the Q1 lenses it is possible to find new values of scaling such that the associated empirical reliability plot matches the identity line: such values are reported in Table 4. These new values produce uncertainties that are systematically bigger than the ones we reported in Sect. 6.3, with minimum median percent increases of around $41\%$ for the Einstein radii uncertainties, up to $136\%$ for the effective radii uncertainties.
Appendix D Uncalibrated aleatoric and epistemic uncertainty trends
Figure 20: Trends of aleatoric (blue) and epistemic (orange) relative uncertainties as a function of the true value of the lens parameters. For the position angle and lens magnitude, we consider the absolute uncertainties. For clarity, we plotted only $5000$ random points from the test set.
We report, in Fig. 20 , an alternative version of Fig. 6, where the trends of the (uncalibrated) aleatoric and epistemic components of the total uncertainty are shown separately. The fact that the epistemic uncertainty is systematically lower than the aleatoric uncertainty is a good sign, because it indicates that, at least for the mock Euclid-like images that compose the test-set, the network is confident that the learned mapping between images and parameters is well determined (i.e. the error associated with a lack of training is irreducible, and the dominant uncertainty is the aleatoric one). This means that more training data of the same type or a larger network is not necessary, because the model has seen enough similar examples to have consistent predictions in the range covered by the training set, and the model is considered expressive enough.
Appendix E Analysis of signal-to-noise ratio of mock Euclid lenses
Figure 21: Total calibrated uncertainty for each lens against total S/N. The median trends of the scatter plots, along with the respective scatters associated with the 16th and 84th percentiles, are shown as red points and bars, respectively.
Figure 22: Same as Fig. 21, but for the discrepancy between predicted and true parameter value for each lens.
[…]
A similar trend is seen in Fig. 22, but on the scatter of the discrepancy-$\mathrm{S/N}$ relation: quantities associated with the mass have a larger discrepancy scatter between predictions and true values at low values of S/N, due to the fact that the source configurations may be harder to detect if the source is not visible enough, while quantities associated with the light have larger discrepancy scatter at high values of S/N, due to the fact that light from the source interferes with the inference from the lens light.